OpenStudy (anonymous):

Solve -2log8(x + 1) = -8

OpenStudy (australopithecus):

isolate the logrithm, so divide both sides by -2, then make both sides of the equation exponents to the number 8 to cancel the logarithm

OpenStudy (anonymous):

can you show me step by step ?please

OpenStudy (australopithecus):

-2log8(x + 1) = -8 = log_8(x+1) = -8/-2 = \[8^{\log_8(x+1)}=8^{-8/-2} \] remember the rule: \[y^{\log_y(x)} = x\] another example of this rule: \[3^{\log_3(x+1)} = x+1\] etc

OpenStudy (australopithecus):

so you end up with x+1 = 8^(8/2)

OpenStudy (australopithecus):

now you can use simple algebra to solve

OpenStudy (australopithecus):

just remember that rule, and the fact that as long as you do the same operation to both sides of the equation the equation maintains its equality

OpenStudy (anonymous):

thank you :)

OpenStudy (australopithecus):

no problem any other questions feel free to ask

OpenStudy (anonymous):

just one more Write 2log35 + log32 as a single logarithm ?

OpenStudy (australopithecus):

use the rule \[xlog(3) = \log(3^{x})\] an example of this rule: \[3\log(x+1) = \log((x+1)^{3})\] and the rules (remember you can only do this with logarithms that are the same) \[\log(x) + \log(y) = \log(xy)\] and \[\log(x) - \log(y) = \log(\frac{x}{y})\] you cannot do this though \[\log_8(x) + \log(y) \neq \log(xy)\]

OpenStudy (australopithecus):

do you follow? an example of the last two rules is \[\log(s) - \log(y) + \log(3x) - \log(m)= \log(\frac{3xs}{ym})\]

OpenStudy (anonymous):

2log(base3)(7) is this right

OpenStudy (australopithecus):

I think it may just be log(37) \[\log(x) = \log_{10}(x)\]

OpenStudy (australopithecus):

people are lazy so we just call log(x) base 10

OpenStudy (australopithecus):

unless the the 3 is in the subscript then it is \[\log_3(x)\]

OpenStudy (australopithecus):

I'm assuming \[ 2\log_{10}35 + \log_{10}(32)\]

OpenStudy (australopithecus):

so the first step \[\log(35^{2}) + \log(32)\]

OpenStudy (anonymous):

2log3^5 + log3^2 this the question sorry

OpenStudy (australopithecus):

\[2\log(3^{5}) + \log(3^{2})\] so this is the question

OpenStudy (australopithecus):

?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

here is the first step just use the rules I showed you I will tell you if you get the right answer \[\log(3^{5(2)}) + \log(3^{2})\]